There are three consecutive integers. If the square of the last plus the product of the other two is 137, find the sum of the integers. If the numbers were three rational number, how would this affect the result?
Answer by mananth(16949)

let the integers be (a-1),a, a+1
a(a-1)+(a+1)^2=137
a^2-a+a^2+2a+1=137
2a^2+a-136=0

2a^2+17a-16a-136=0
2a(a+8.5)-16(a+8.5)=0
(2a-16)(a+8.5)=0
2a=16 a=8
7,8,9
24 is the sum

OR a=-8.5
not an integer so ignore